# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 7 Conclusions
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\frac{1+e^{2t}-\mathrm{i}\,}{1+e^{2t}+\mathrm{i}\,}\,,\qquad u_{2}^{\prime}=-\frac{2\,\mathrm{i}\,e^{t}}{1+e^{2t}+\mathrm{i}\,}\,$, (A.3)
so that $|u_{1}^{\prime}|=1$ and $u_{2}^{\prime}$ represents a bright soliton. This solution has a special structure, see Appendix Appendix B: An almost trivial class of solutions of the two-component vector FL equation.
@@
Figure 14: Plots of the absolute squares of the two components, and of $\|u^{\prime}\|^{2}$, for the solution (A.1) of the two-component vector FL equation. Comparison with Fig. 15 shows that the dark, respectively bright soliton part for $t>0$ appears since $c_{2}\neq 0$. Correspondingly, the bright soliton part of $u_{2}^{\prime}$ for $x>0$ shows up because $c_{1}\neq 0$.
Figure 15: Plots of the absolute squares of the two components of the solution (A.2). The data are the same as used in Fig. 14, except that now $c_{1}=0$.
@@
Appendix B: An almost trivial class of solutions of the two-component vector FL equation
@@
with complex constants $A_{1},A_{2}$. (A.3) has the form of our ansatz, but $|u_{2}^{\prime}|$ is not constant. This apparent contradiction is resolved by noting that, in this example, we have $\alpha_{2}=0$ and $|f_{1}|^{2}=\alpha_{1}^{-1}$, so that only $|f_{1}|$ is conserved and left and right hand sides of the ODE for $f_{2}$ vanish separately, hence $f_{2}$ can be a any function of $t$.
Appendix C Superposition of two solutions from the class in Example 6.2

### v2

@@ 7 Conclusions
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\frac{1+e^{2t}-\mathrm{i}}{1+e^{2t}+\mathrm{i}}\,,\qquad u_{2}^{\prime}=-\frac{2\,\mathrm{i}\,e^{t}}{1+e^{2t}+\mathrm{i}}\,$, (B.3)
so that $|u_{1}^{\prime}|=1$ and $u_{2}^{\prime}$ represents a bright soliton. This solution has a special structure, see Appendix Appendix C: An almost trivial class of solutions of the two-component vector FL equation.
@@
Figure 16: Plots of the absolute squares of the two components, and of $\|u^{\prime}\|^{2}$, for the solution (B.1) of the two-component vector FL equation. Comparison with Fig. 17 shows that the dark, respectively bright soliton part for $t>0$ appears since $c_{2}\neq 0$. Correspondingly, the bright soliton part of $u_{2}^{\prime}$ for $x>0$ shows up because $c_{1}\neq 0$.
Figure 17: Plots of the absolute squares of the two components of the solution (B.2). The data are the same as used in Fig. 16, except that now $c_{1}=0$.
@@
Appendix C: An almost trivial class of solutions of the two-component vector FL equation
@@
with complex constants $A_{1},A_{2}$. (B.3) has the form of our ansatz, but $|u_{2}^{\prime}|$ is not constant. This apparent contradiction is resolved by noting that, in this example, we have $\alpha_{2}=0$ and $|f_{1}|^{2}=\alpha_{1}^{-1}$, so that only $|f_{1}|$ is conserved and left and right hand sides of the ODE for $f_{2}$ vanish separately, hence $f_{2}$ can be any function of $t$.
Appendix D: Superposition of two solutions from the class in Example 6.5